) What is the oxidation number of cobalt in Co2(SO4)3?
Answer #1

oxidation number of cobalt is : +3

Explanation

Sulfate ion SO42- has -2 oxidation number

Total sulfate ions in Co2(SO4)3 are 3

Total oxidation number of sulfate ions = 3 * (-2) = -6

To maintain electrical neutrality, total oxidation number of Co = -(Total oxidation number of sulfate ions)

total oxidation number of Co = -(-6)

total oxidation number of Co = 6

oxidation number of Co = (total oxidation number of Co) / (total Co atoms)

oxidation number of Co = (6) / (2)

oxidation number of Co = +3